### v1

# E-TRIALS: Empowering Data-Driven Decisions to Enhance Computer-Based Learning Platforms

Abstract
Computer-based learning platforms (CBLPs) have become a common medium in schools, transferring how students learn and interact with education content. A recent report by the Institute of Education Sciences revealed that 94% of public schools now provide a computer for every student. The high percentage of students having their own computers highlights the need for researchers to investigate how students learn in CBLPs and identify instructional designs that can enhance their performance. In this paper, we introduce E-TRIALS, a tool by ASSISTments designed to conduct experiments and analyze student performance. Using data from three experiments, we propose a generalizable approach for analyzing experimental data and modeling results. We evaluate three Average Treatment Effect (ATE) estimators: Student’s t-test $\hat{\tau}_{t-test}$, Ordinary Least Squares regression $\hat{\tau}_{OLS}$, and Leave-One-Out Potential outcomes (LOOP) $\hat{\tau}_{LOOP}$. Our approach can be used for future E-TRIALS experiments. Code used for this work is available at https://osf.io/xp6ch/.
Keywords: Educational Interventions, E-TRIALS, A/B Testing, Learning Analytics
1 Introduction
In 1923, Thomas Edison predicted that motion pictures would revolutionize the education system within 20 years by replacing teachers and textbooks entirely [Salisbury (1986)]. This prediction reflects a recurring pattern in the education system, which is the introduction of new technologies often sparks enthusiasm and bold claims about their potential to transform learning. For instance, in 1961, the Ford Foundation invested over $10$ million in a program to promote televised instruction. The foundation believed televised instruction could ”equal or surpass conventional methods” and make superior teaching resources accessible to more students [Ford Foundation (1961)]. Some of the funds were even used to release teachers from their daily duties to serve as ”television consultants.”
This optimism was not unique to television. Similar enthusiasm emerged with the widespread adoption of computers in education. Educators quickly embraced computers for administrative tasks, research, and instructional purposes [Billings and Moursund (1988)]. In the 1960s, the National Science Foundation spearheaded efforts to integrate computers into classrooms aiming to enhance instruction. At the time, only 1% of secondary schools had access to computers, but by 1975, this figure had risen to 23% [Com (1997)]. Initiatives like the ”Computer Literacy Curriculum” were introduced to ensure students and teachers were adequately prepared for this technological shift [Molnar (1975)].
The growing presence of computers in schools paved the way for intelligent tutoring systems (ITS), which aimed to deliver personalized instruction, provide immediate feedback, and assess student knowledge [Billings and Moursund (1988)]. One of the earliest examples, SOPHIE (SOPHisticated Instructional Environment), was developed in 1975 by Brown et al. SOPHIE simulated electronic equipment and allowed students to practice troubleshooting without the risk of damaging real devices [Brown et al. (1974)]. Leveraging techniques from artificial intelligence, cognitive science, and psychology, SOPHIE was able to understand natural language inputs and respond in a supportive tone within two seconds. Additionally, SOPHIE outperformed human observers in identifying faults and was more effective than students at diagnosing errors in their reasoning.
Another notable advancement in ITS was the development of cognitive tutors based on the Adaptive Control of Thought (ACT) theory of learning and problem-solving [Anderson et al. (1995)]. These tutors were guided by eight principles, such as representing student competence as a set of production rules, providing immediate feedback, and minimizing cognitive load. Researchers applied these principles to create tutors for subjects like LISP, geometry, and algebra. The LISP tutor helped students complete exercises 30% faster and achieve 43% higher scores on a post test. The algebra tutor initially didn’t show significant results, but a later version showed significant achievement gains. The geometry tutor helped improve student’s grades by more than one letter grade on a test.
Over the years, terms like intelligent tutoring systems (ITS), computer-assisted instruction (CAI), and online learning platforms have often been used interchangeably. While these systems share similarities, such as providing immediate feedback and facilitating learning, they also have distinct features. For instance, ITS are typically knowledge-centered, whereas CAI tends to be interaction-centered [Larkin and Chabay (2021)]. To avoid confusion, this paper uses the term computer-based learning platforms (CBLPs) to refer to systems that: 1) allow students to interact with simulations or problem-solving tasks, 2) provide personalized or generic feedback, 3) adapt instructional content or assessments to the student’s knowledge level, 4) incorporate multimedia elements (e.g., text, audio, graphics, video), and 5) track student progress and generate performance reports for students and teachers.
The adoption of CBLPs in classrooms has surged in recent years, largely due to the increase of availability of computers to each student. Before the COVID-19 pandemic, 45% of public schools reported having a computer for every student [Gray and Lewis (2021)]. By the 2022-2023 academic year, this figure had rise to 94% [Institute of Education Sciences (nd)]. Unlike earlier technologies, CBLPs have consistently demonstrated their ability to enhance learning outcomes, particularly when designed to support rather than replace teachers [Anderson (2014), Murphy et al. (2020), Leelawong and Biswas (2008), Toyama (2011)]. These platforms have transformed how students learn, complete assignments, and receive assessments. For example, CBLPs enable students to receive immediate and personalized feedback, while teachers benefit from analytical dashboards that provide insights into student performance. These tools allow educators to quickly identify struggling students and tailor their support accordingly.
Despite their advantages, developing effective CBLPs is a complex challenge due to the vast array of instructional design choices available. As Koedinger et al., noted in Science, there are approximately 300 trillion possible design configurations [Koedinger et al. (2013)]. Researchers are struggling to identify which instructional choices to include and omit. Instructional choices, such as how often should students receive feedback, and what type of feedback is most effective? Should high-performing students receive delayed feedback, while low-performing students receive immediate feedback? Should instruction rely solely on text, or should it incorporate multimedia elements like videos, graphics, or audio? Are open-ended questions more effective than multiple-choice questions? These are just a few of the many decisions researchers and developers must navigate.
The overwhelming number of possible design choices birthed the term ”the rise of the super experiment” [Stamper et al. (2012)]. This concept refers to data-driven educational experiments conducted at the internet level to identify which instructional choices are most effective and to whom. To address this need, tools like UpGrade by Carnegie Learning and TerraCotta within the Canvas Learning Management System have been developed to facilitate such experiments [Fancsali et al. (2022)]. However, many of these tools are either closed-source or lack the flexibility to explore diverse research questions across various aspects of CBLPs.
In this paper, we introduce Ed-Tech Research Infrastructure to Advance Learning Sciences (E-TRIALS) a free and open tool developed by ASSISTments. E-TRIALS enables researchers to conduct experiments in any subject area. We describe its features, the types of experiments it supports, and how it can address critical research questions in learning sciences and technologies. Additionally, we highlight how E-TRIALS simplifies data analysis by providing consistent and similar structure of the experiment data. To demonstrate this capability, we selected three experiments in E-TRIALS and used the same code to analyze and report findings. To determine the effectiveness of the experiments, we used three Average Treatment Effect (ATE) estimators: Student’s t-test $\hat{\tau}_{t-test}$, Ordinary Least Squares regression $\hat{\tau}_{OLS}$, and Leave-One-Out Potential outcomes (LOOP) $\hat{\tau}_{LOOP}$.
2 Experimental Environment
ASSISTments is a web-based and free-to-use CBLP designed to assist teachers, rather than replace them [Heffernan and Heffernan (2014)]. The platform allows teachers to assign mathematical problems from open-source curriculum, such as Illustrative Mathematics, or to create their own problem sets. As students work through problem sets, they can request support, which would appear in the form of hints or explanations. ASSISTments is widely used, with over 500,000 students and 30,000 teachers annually [ASSISTments (2025)].
2.1 E-TRIALS
The Ed-Tech Research Infrastructure to Advance Learning Sciences (E-TRIALS) is a built-in research functionality within ASSISTments. It enables researchers to conduct experiments in the realm of learning sciences, without any coding expertise. Before the wide spread of CBLP and tools like E-TRIALS, researchers relied on manual methods to design and conduct experiments [Angrist (2003)]. Researchers would rely on a coin toss to assign students to a condition. Fortunately, E-TRIALS simplifies this process of conducting experiments and provides researchers with in-depth data about the student. With over 500,000 active student users and more than 30,000 teacher users [ASSISTments (2025)], ASSISTments offers researchers access to diverse study participants from various background.
2.1.1 Previous Experiments in E-TRIALS
E-TRIALS is one of five digital platforms funded by Institute of Education Sciences (IES) to develop infrastructure supporting research in the field of learning sciences and technologies [E-TRIALS (2025)]. E-TRIALS has already facilitated impactful studies, such as work by [Walkington et al. (2019)], who used E-TRIALS to explore how language skills can impact student performance in algebra and evaluate their effects on accuracy and response time. In the study, they manipulated six language features: 1) the number of sentences, 2) pronouns, 3) word concreteness, 4) word hypernymy, 5) sentence consistency, and 6) problem topic. The study had 451 students and 2,825 observations. Each observation represented a student’s response to a problem. While the findings showed little evidence that language features significantly impact mathematics word problem-solving performance, the findings did reveal that sentence consistency reduce response time.
Another notable study that used E-TRIALS was work by [Vanacore et al. (2023)]. They evaluated the impact of non-cognitive interventions on student learning and behavior. This study included five experiments: 1) embracing mistakes, 2) inspirational quotes, 3) social comparison, 4) emotion labeling, and 5) confidence judgments. The experiments had 20,472 students. Among the five experiments, only social comparison produced a positive effect on post-test performance.
In a broader analysis, [Prihar et al. (2022)] evaluated 50 different experiments conducted using E-TRIALS. These experiments investigated a wide range of research questions, such as the effectiveness of explanation vs. answer-only feedback, video examples vs. text-based, common wrong answer feedback vs. no feedback, fill-in vs. multiple-choice questions, and motivational vs. non-motivation messages. Collectively, these experiments had a 30,408 students. The primary goal of this meta-analysis was to identify consistent patterns across experiments, such as personalization and effective tutoring methods within CBLPs.
These examples are an indication of the types of researcher questions E-TRIALS can answer, and the rich data that researchers receive. E-TRIALS opens new avenues for addressing novel research questions and generating data-driven insights. Its ability to handle large sample sizes, diverse interventions, and complex experimental designs makes it an invaluable tool for advancing research.
2.1.2 Developing an Experiment
Experiments in E-TRIALS are delivered through ASSISTments. They can be assigned using learning management systems, such as Google Classroom or Canvas. As show in Figure 1, E-TRIALS supports three types of experiments: single support, support comparison, and problem varied.
Figure 1: The 3 Types of Studies in E-TRIALS
Single Support: Single support studies are designed to evaluate the effects of a specific type of student support message, such as a hint or explanation. These messages can be in the form of text, images, videos, or audio. The control condition in these studies is the best-so-far condition, which is ASSISTments’ best selection of a support to deliver to maximize the student’s performance.
Researchers use ASSISTments’ math problem content library, which comprises problems from open-source curricula, such as Kendall Hunt, Illustrative Mathematics, and Engage NY/Eureka Math to create experimental conditions. Researchers do not need to recruit students because they’re randomly assigned to a condition.
Support Comparison: Support comparison studies build on the single support model by allowing researchers to investigate two or more types of student support messages. Examples include comparing hints vs. explanations, hints vs. worked examples, or explanations vs. a dynamic algebraic notation system. As with single support studies, researchers can use ASSISTments’ existing math content to create conditions. For example, [Smith et al. (2022)] used a support comparison study to explore the impact of worked example extensiveness and the degree of dynamic presentation on student learning. Their study had six conditions: 1) static concise, 2) static extended, 3) sequential concise, 4) sequential extended, 5) dynamic history, and 6) dynamic no history. The results revealed an overall improvement from pretest to post test after viewing the worked examples, but there was no significant differences on post test performance between worked example presentation.
Problem Varied: Problem varied studies offer researchers the flexibility to create their own problem content and recruit participants. These studies accommodate a broader range of participants, from K-12 to higher education students. [Smalenberger and Smalenberger (2022)] conducted a problem varied study to examine how middle and college students benefited from three types of problem support messages: text hints, explanations, and video explanations. They compared the effects of these formats within the college sample, and between the college and middle school samples. Findings revealed that for college students, on-demand video explanations was more effective than text hints. Problem varied studies are not limited to math-related research questions. For example, [Unal et al. (2020)] used this study type to design a think-aloud study that explored cognitive control and rule learning in a real-world setting. Their study featured probability problem sets and recruited undergraduate students as participants.
2.1.3 Deploying an Experiment
E-TRIALS supports two deployment methods depending on the study type: automatic deployment and self-deployment.
  • Automatic Deployment: Studies categorized as single support or support comparison are automatically deployed. These studies are added to ASSISTments’ content library, where teachers can assign them to students. Randomization occurs at the student level within the platform.
  • Self-Deployment: Problem varied studies require self-deployment because researchers need to recruit their own participation pool. Researchers collaborate with students, teachers, or schools to assign content. Randomization for these studies can occur at various levels, include student, teacher, class, or school.
2.1.4 Completion of Experiment
Once a study has been deployed, researchers can request the study data after reaching a specific number of student participants or after a predetermined time frame. The choice between these options depends on the study’s requirements.
Researchers will receive eight distinct data files, each containing specific insights into student interactions, learning outcomes, and experimental conditions. These files provide a comprehensive view of the study’s results, ranging from click stream data to problem difficulty and demographic context.
Below is a detailed description of each data file:
  1. 1.
    Action Logs: This file captures students’ click stream data during an assignment. It records every action a student takes, such as starting, resuming, or finishing a problem, requesting a support message or submitting an answer. Each row represents a single student action.
  2. 2.
    Problem Logs: This file contains problem-level data, such as time spent on the problem, the number of attempts, and scores. ASSISTments uses two scoring metrics:
    • Discrete Score: Either 0 or 1.
    • Continuous Score: Values can be 0, 0.33, 0.67, or 1.0. Students lose 0.33 points each time they request a hint. Meaning, the highest possible score a student can receive after viewing one hint is 0.67.
  3. 3.
    Assignment Logs: This file summarizes assignment-level data. Each row includes information on total time spent on the assignment, the total number of attempts or hints, and the cumulative discrete score for the assignment.
  4. 4.
    Prior Logs: This files provides class- and school-level data related to students’ prior learning. For the class-level data, it includes the curriculum used, the class grade level, the number of students in the class, and the class’s average performance. For the school-level data, we get this data from the National Center for Education Statistics (NCES) [Hussar and Bailey (2011)] based on the teacher’s school address. It includes demographic data, such as percentage of students receiving free or reduced lunch, ethnicity and gender distribution, and the school’s location (rural, urban, or suburban). The school_id provided in this file can be used to access additional school-level, or even district-level data through NCES or the Education Data Explorer by the Urban Institute. If a teacher does not provide their school address, however, it is not feasible to link the data for their class to school-level data.
  5. 5.
    Same Skill Prior Logs: This file is similar to the Prior Logs file. However, this file focuses only on prior problems by the student that share the same skill codes as those in the experiment.
  6. 6.
    Same Skill Post Logs: This file is similar to the Same Skill Prior Logs, but it pertains to problems completed after the experiment.
  7. 7.
    Assignment Settings: The file details the settings applied to the assignment. Assignment settings be setting a time limit or enabling the redo feature. The redo feature allows students to attempt a similar, but not the same problem again if a student fails to achieve a perfect score on the initial problem.
  8. 8.
    Redo Logs: If the redo feature is enabled, this file provides problem-level data for the redo attempts, which shows how students performed on these additional problems.
These eight files are available only for studies conducted with participants from ASSISTments. Studies involving external participants (e.g. college students recruited for problem-varied studies) will not include data on students’ prior and post performance in ASSISTments. To view an example of the eight data files, click here.
3 Example Studies
3.1 The Three Experiments
One unique advantage of E-TRIALS is the consistent format and structure of the eight data files provided to researchers after the completion of an experiment. This consistency can help researchers reuse the same code across multiple experiments. To demonstrate this advantage, we selected three support comparison studies within E-TRIALS and applied the same code to evaluate all three experiments. A link to the code is provided in the Data Cleaning & Preparation section.
The three experiments are summarized below:
  1. 1.
    Experiment 14 (similar vs. different example feedback): This experiment explored math learning and persistence by comparing two feedback conditions:
    • Similar Example Feedback: Feedback that aligns closely with the problem at hand.
    • Different Example Feedback: Feedback that diverges from the problem’s context.
    Students were randomly assigned to one of the two conditions. Figure 2 illustrates the differences between these feedback types.
    Figure 2: The support messages students received in both conditions
  2. 2.
    Experiment 15 (drawing vs. imagining the problem vs. image provided): This study investigate the impact of visualization techniques. Students were randomly assigned to one of three conditions:
    • Drawing on Paper: Students were asked to physically draw the image in the problem.
    • Mental Visualization: Students visualized the image mentally without drawing.
    • Provided Drawing: Students were given a completed drawing of the image.
    Figure 3 highlights the differences between the three conditions.
    Figure 3: The support messages students received in each condition
  3. 3.
    Experiment 16 (reflect vs. regular hint): This experiment examined the effectiveness of encouraging students to reflect on their confusion before requesting an additional hint.
    • Reflect: Students were prompted to reflect on their confusion before requesting additional hints.
    • Control: Students received hints as usual without any prompts to reflect.
    Students were randomly assigned to one of two conditions. Figure 4 displays the reflection prompt.
    Figure 4: The support messages students received in both conditions
3.2 Data Cleaning & Preparation
In this section, we outline the steps taken to clean and prepare the data. The code used for this process can be accessed here.
The data cleaning process began with the Problem Logs table. Table 1 shows a summary of the conditions, student and assignment scores, their means and standard deviations at the condition level, and the average number of problems assigned per condition.
Table 1: Experiments Initial Summary
Experiments Condition Condition Type Student Count Assignment Count Mean Std. Dev. Avg. Problems Assigned
14 0 Similar 343 362 0.565 0.448 10.06
1 Different 348 365 0.573 0.445 9.81
2 Best-So-Far 334 362 0.571 0.446 10.54
15 0 Drawing 97 113 0.747 0.393 18.33
1 Imagining 98 109 0.679 0.432 18.02
2 Image Provided 112 133 0.758 0.388 19.41
3 Best-So-Far 111 137 0.779 0.378 22.36
16 0 Reflect 517 582 0.702 0.418 10.62
1 Control 488 552 0.708 0.412 10.32
2 Best-So-Far 516 593 0.709 0.412 10.38
To ensure we analyze the experiments correctly, several filtering steps were applied. First, we removed ASSISTments’ default condition, best-so-far because it provides students with different support messages using different format. Meaning, one student could receive a support message in a text format, whereas another student could receive it in a video. Due to the inconsistency, we decided to remove the condition from our analysis.
Next, we removed students who did not request at least one hint during the assignment, as hint interaction was required to determine the effect of the experiment. This filtering step resulted in the removal of 400 students in experiment 14, 109 students in experiment 15, and 477 students in experiment 16. Additionally, sub-problems were treated as independent problems. The updated summary of the dataset after applying these filters is presented in Table 2.
To assess the extend to student participation, we calculated the attrition rate. We calculated the attrition rate by using the proportion of missing values in the post and prior continuous score variables that’s in the Same Skill Prior & Post Logs tables relative to the total number of observations [Schafer and Graham (2002)]. The attrition rates for the three experiments were significantly low, indicating that the vast majority of participants competed the study.
Table 2: Experiments Summary After Filtering & Cleaning
Experiments Condition Condition Type Student Count Assignment Count Mean Std. Dev. Attrition Rate
14 0 Similar 157 161 0.427 0.454 0.0047
1 Different 134 137 0.396 0.448 0.0035
15 0 Drawing 64 66 0.649 0.439 0.0015
1 Imagining 66 74 0.628 0.449 0.0019
2 Image Provided 68 75 0.682 0.424 0.002
16 0 Reflect 282 296 0.574 0.458 0.00004
1 Control 246 260 0.554 0.455 0.00005
Next, the filtered Problem Logs table was joined with the Assignment Logs table to produce an assignment-level dataset. This table included details about assignment and teacher IDs, and was further aggregated to calculate the number of problems assigned to each student and the number they completed.
3.3 Feature Selection in Covariate Adjustment
It is essential for students to have comparable prior knowledge and learning motivation. To minimize the influence of confounding factors, we aim to enhance comparability through covariate adjustment. Fortunately, our dataset contains a large number of covariates, which we leverage to account for variation in students’ baseline knowledge.
First, we utilized students’ prior and post continuous scores from assignments associated with the same skill code as the experimental problems. This data was drawn from the Same Skill Prior Logs and Same Skill Post Logs. Second, we analyzed the Prior Logs table to identify covariates potentially correlated with student performance. Although numerous covariates are available, feature selection is necessary to ensure computational efficiency, particularly for OLS regression. For Random Forest models, however, feature selection is handled inherently, as the algorithm randomly selects subsets of features during the splitting process.
We began by subjectively selecting a set of covariates. Next, we calculated the linear correlations between these covariates and student performance, visualizing the results in Figure 5. While focusing solely on linear correlation is not the most comprehensive approach, as it overlooks non-linear relationships, we chose this measure due to its simplicity, ease of application, and suitability for OLS regression. [^1] OLS regression and LOOP, based on Random Forest, will be discussed in detail in the next section.
Figure 5: Feature Selection
Based on Figure 5, we excluded covariates in the plot that appeared close to white, as they indicated weak correlations. This process allowed us to identify covariates with meaningful correlations to the outcomes through a data-driven approach.
Finally, the assignment-level dataset was combined with the prior and post logs to create the final dataset used for analysis.
4 Method
In this section, we discuss the statistical methods employed to assess the impact of different types of tutoring support on students’ proximal and distal performance. We also provide the rationale for selecting these methods, emphasizing their respective advantages and limitations. The methods are presented in formal mathematical terms for clarity and rigor. The two outcome variables we explored are distal and proximal learning. Distal learning refers to the knowledge students retain after the experiment. It is measured by the average score students received on problem sets post-experiment, which assess the same skills covered during the experiment. Proximal learning refers to the knowledge students gain during the experiment. It is measured by the average continuous score students received during the experiment problem set.
Consider an experiment with $N$ participants, indexed as $i=1,2,...,N$. Each participant can be assigned to either treatment or control arm at random, and we denote $T_{i}$ following Bernoulli distribution be the $i$ th participant’s treatment assignment, that is,
$T_{i}$ $\perp T_{j},\quad i\neq j$ (1)
$P(T_{i}=t)$ $=p_{i}^{t}(1-p_{i})^{1-t},\quad t\in\{0,1\}$
where $T_{i}=1$ when the $i$ th participant is assigned to treatment arm and $T_{i}=0$ when control arm. Furthermore, we observe outcomes $Y_{i}$ and $q$-dimensional baseline covariates $Z_{i}$ besides treatment assignment $T_{i}$ for each participant.
Assume each participant has two nonrandom potential outcomes– treatment potential outcomes $t_{i}$ and control potential outcomes $c_{i}$. We can write observed outcomes $Y_{i}$ as
$Y_{i}=T_{i}t_{i}+(1-T_{i})c_{i}$ (2)
Individual Treatment Effect (ITE) $\tau_{i}$ can be written as
$\tau_{i}=t_{i}-c_{i}$ (3)
and the Average Treatment Effect (ATE) $\bar{\tau}$ as
$\bar{\tau}=\frac{1}{N}\sum_{i=1}^{N}\tau_{i}$ (4)
which serves as our vital interested parameter to be estimated.
Below we introduce LOOP estimator $\hat{\tau}_{LOOP}$ along with the other two common ATE estimators: Student’s t-test estimator $\hat{\tau}_{t-test}$ and Ordinary Least Squares (OLS) regression estimator $\hat{\tau}_{OLS}$. We also explore their strengths and limitations.
4.1 Student’s t-test
$\hat{\tau}_{t-test}$ represents the difference of the average of the observed treatment outcomes and the average of the observed control outcomes.
$\hat{\tau}_{t-test}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}-\frac{1}{N-n}\sum_{i=1}^{N-n}Y_{i}$ (5)
where $n$ be the number of participants in treatment arm. It provides an unbiased estimate of the average treatment effect with $0<n<N$.
4.2 OLS Regression
Outcomes $Y_{i}$ and $q$-dimensional baseline covariates $Z_{i}$ are available in E-Trials, enabling the implementation of covariate adjustments for estimating the ATE. OLS regression estimators $\hat{\tau}_{OLS}$ naturally arise when considering covariate adjustment.
$Y_{i}=\tau_{OLS}T_{i}+\beta X_{i}+\epsilon_{i}$ (6)
where $\hat{\tau}_{OLS}$ can serve as one estimator for ATE $\tau$.
The estimator $\hat{\tau}_{\text{OLS}}$ leverages information from both treatment assignment $T_{i}$ and $q$-dimensional baseline covariates $Z_{i}$, providing a more precise estimate compared to $\hat{\tau}_{\text{t-test}}$. However, $\hat{\tau}_{\text{OLS}}$ has two notable limitations: (a) it may be biased since the assumption $X_{i}\perp T_{i}$ is violated, and (b) it fails to account for potential non-linear relationships between $Y_{i}$ and $X_{i}$.
4.3 LOOP
To address the limitations of the OLS estimator, we propose using LOOP to estimate $\tau$. $\hat{\tau}_{LOOP}$ offers an unbiased estimate of ATE while incorporating covariate adjustments. We present the steps of the LOOP methodology. A key quantity, $m_{i}$, is defined for each participant as follows:
$m_{i}=(1-p_{i})t_{i}+p_{i}c_{i}$ (7)
and a (signed) inverse probability weight $U_{i}$, defined as follows:
$U_{i}=\begin{cases}\frac{1}{p_{i}},&T_{i}=1,\\ -\frac{1}{1-p_{i}},&T_{i}=0.\end{cases}$ (8)
and note that $U_{i}$ has expectation 00.
Now, consider $\hat{\tau}_{i}=(Y_{i}-\hat{m}_{i})U_{i}$ as an unbiased estimator of ITE.
When $\hat{m}_{i}\perp U_{i}$, $\hat{\tau}_{i}$ can serve as an unbiased estimator of ITE. This can be shown as follows:
$E(\hat{\tau}_{i})$ $=E[(Y_{i}-\hat{m}_{i})U_{i}]$ (9)
$=E(Y_{i}U_{i})-E(\hat{m}_{i})E(U_{i})$
$=\tau_{i}$
Additionally, we can demonstrate that $\hat{\tau}_{i}$ exhibiting reduced variance, provided that $\hat{m}_{i}$ closely approximates $m_{i}$. The variances can be calculated as follows:
$Var(\hat{\tau}_{i})=\frac{1}{p_{i}(1-p_{i})}E[(\hat{m}_{i}-m_{i})^{2}]$ (10)
To summarize, $\hat{\tau}_{i}$ is not only unbiased but also exhibits reduced variance when (a) $\hat{m}_{i}\perp T_{i}$ and (b) $\hat{m}_{i}$ serves as an accurate estimator of $m_{i}$.
We now introduce Leave-One-Out imputation. Let the LOOP estimator of ATE, $\hat{\tau}_{LOOP}$, be denoted as:
$\hat{\tau}_{LOOP}=\frac{1}{N}\sum_{i=1}^{N}\hat{\tau}_{i}$ (11)
where $\hat{\tau}_{i}=(Y_{i}-\hat{m}_{i})U_{i}$, and $\hat{m}_{i}$ is obtained using Leave-One-Out imputation, i.e., we sequentially exclude each observation $i$ and use the remaining $N-1$ observations to impute $t_{i}$ and $c_{i}$. The imputation method is flexible and can be customized as needed. We define $\hat{m}_{i}$ as follows:
$\hat{m}_{i}=(1-p_{i})\hat{t}_{i}+p_{i}\hat{c}_{i}$ (12)
By leaving out the $i$-th observation when imputing $\hat{m}_{i}$, we ensure that $\hat{m}_{i}\perp T_{i}$, guaranteeing that $\hat{\tau}_{i}$ remains unbiased. Consequently, $\hat{\tau}_{LOOP}$ is also unbiased, regardless of the algorithm used to impute $t_{i}$ and $c_{i}$. Note that both $\hat{t}_{i}$ and $\hat{c}_{i}$ are imputed, even when one is observed, to maintain the independence $\hat{m}_{i}\perp T_{i}$.
When it comes to methods for imputing potential outcomes, simply ignoring covariates during imputation makes the LOOP estimator equivalent to the student’s t-test estimator. However, the LOOP estimator outperforms $\hat{\tau}_{\text{SD}}$ as long as the imputation of potential outcomes improves beyond the baseline mean imputation approach. In the work by Wu and Gagnon-Bartsch [Wu and Gagnon-Bartsch (2018)], Random Forest (RF) is recommended for imputing $t_{i}$ and $c_{i}$. This is because RF, as an enhancement over individual decision trees, not only improves accuracy but also provides automatic variable selection, resulting in more computational efficiency.
In this paper, we calculate the t-test estimator $\hat{\tau}_{\text{t-test}}$, the OLS regression estimator $\hat{\tau}_{\text{OLS}}$, and the LOOP estimator $\hat{\tau}_{\text{LOOP}}$ to demonstrate that the LOOP estimator is an unbiased estimator of the ATE and achieves greater precision by adjusting for baseline covariates. Furthermore, when using the LOOP estimator with Random Forest (RF), we can identify non-parametric relationships within the data that the OLS estimator fails to capture.
5 Results & Discussions
In this section, we present the results of the example experiments. Tables 3, 3, and 4 summarize the ATE estimates, $\hat{\tau}_{t-test}$, $\hat{\tau}_{OLS}$, and $\hat{\tau}_{LOOP}$, along with their respective variation and significance measures .
First, we present the results of the analysis for Experiment 14, shown as Table 3, which aims to determine whether providing students with feedback that either aligns closely with the problem at hand or diverges from its context influences their proximal and distal-term performance. We include the estimates, standard errors, and $p_{value}$ of the results, accompanied by a figure visualizing the confidence intervals. This visualization helps us explore whether there is any likely significant positive or negative effect.
Table 3: Experiment 14: Similar vs. Different Example Feedback
Performance Method Estimate Std. Error $\mathbf{Pr(>|t|)}$
Proximal $\hat{\tau}_{t-test}$ 0.0254 0.0333 0.4461
$\hat{\tau}_{OLS}$ 0.0233 0.0290 0.4240
$\hat{\tau}_{LOOP}$ 0.0073 0.0318 0.8201
Distal $\hat{\tau}_{t-test}$0.0053 0.0285 0.8516
$\hat{\tau}_{OLS}$0.0141 0.0199 0.4814
$\hat{\tau}_{LOOP}$0.0219 0.0237 0.3584
  • Note: The smallest Std.Error in each performance group is in bold.
Figure 6: Experiment 14: Confidence Intervals
From Table 3, we observe no significant difference between the similar and different feedback groups. However, after covariate adjustment, the standard errors, as indicated by the Std.Error column, are reduced. The results show that $\hat{\tau}_{OLS}$ achieves the smallest standard error for both proximal and distal performance. It is important to note, though, that $\hat{\tau}_{OLS}$ is biased, whereas $\hat{\tau}_{LOOP}$ is unbiased and exhibits the second smallest standard error. These findings confirm that while both OLS regression and LOOP provide more precise estimators compared to student’s t-test, only LOOP offers an unbiased estimate. Thus, it is reasonable to select LOOP as the preferred estimator for estimating ATE. From Figure 6, we observe that all confidence intervals include $y=0$, indicating that none of the estimates are statistically significant. However, there is a noticeable trend: for proximal performance, students receiving different feedback tend to perform better, whereas for distal performance, they tend to perform worse. This is reflected in the proportion of the spread of the confidence intervals, which leans positive for proximal performance and negative for distal performance. Additionally, we can see that the lengths of the confidence intervals for $\hat{\tau}_{OLS}$ and $\hat{\tau}_{LOOP}$ are shorter than those for the $\hat{\tau}_{t-test}$, suggesting that OLS and LOOP provide more precise estimates compared to the student’s t-test.
Next, we interpret the results of Experiment 15, which examines the impact of different visualization techniques: drawing, mental visualization, and provided images on students’ performance in both proximal and distal terms. The calculated results are presented in Table 4, with corresponding visualizations shown in Figures 7, 8, and 9.
Table 4: Experiment 15: Drawing vs. Imaging vs. Image Provided
Condition Performance Method Estimate Std. Error $\mathbf{Pr(>|t|)}$
Drawing vs. Imaging Proximal $\hat{\tau}_{t-test}$0.0465 0.0352 0.1897
$\hat{\tau}_{OLS}$0.0469 0.0291 0.1091
$\hat{\tau}_{LOOP}$0.0469 0.0291 0.1091
Distal $\hat{\tau}_{t-test}$0.0293 0.0323 0.3651
$\hat{\tau}_{OLS}$0.0352 0.0198 0.0774
$\hat{\tau}_{LOOP}$0.0352 0.0198 0.0774
Imaging vs. Image Provided Proximal $\hat{\tau}_{t-test}$ 0.0472 0.0360 0.1919
$\hat{\tau}_{OLS}$ 0.0377 0.0309 0.2238
$\hat{\tau}_{LOOP}$ 0.0377 0.0309 0.2238
Distal $\hat{\tau}_{t-test}$ 0.0559 0.0330 0.0928
$\hat{\tau}_{OLS}$ 0.0405 0.0189 0.0346
$\hat{\tau}_{LOOP}$ 0.0405 0.0189 0.0346
Drawing vs. Image Provided Proximal $\hat{\tau}_{t-test}$ 0.0008 0.0331 0.9814
$\hat{\tau}_{OLS}$0.0025 0.0297 0.9341
$\hat{\tau}_{LOOP}$0.0025 0.0297 0.9341
Distal $\hat{\tau}_{t-test}$ 0.0266 0.0288 0.3575
$\hat{\tau}_{OLS}$ 0.0196 0.0177 0.2719
$\hat{\tau}_{LOOP}$ 0.0196 0.0177 0.2719
  • Note: The smallest standard error within each performance group and the significant $p$-values are in bold.
Figure 7: Experiment 15: Confidence Intervals
Figure 8: Experiment 15: Confidence Intervals
Figure 9: Experiment 15: Confidence Intervals
In Table 4, we observe that $\hat{\tau}_{OLS}$ and $\hat{\tau}_{LOOP}$ provide the same estimates and standard errors to four decimal places. Both methods exhibit smaller standard errors compared to $\hat{\tau}_{t-test}$. However, it is important to note that $\hat{\tau}_{LOOP}$ is unbiased, making it the recommended estimator. An interesting finding in this experiment is that, when comparing the Imaging vs. Image Provided groups, both the OLS regression and LOOP methods produce significant $p_{value}<0.05$. However, after applying a multiple pairwise comparison adjustment, the significance of these $p_{value}$ disappears. Nonetheless, this result suggests that the two groups may exhibit a meaningful difference if we can further improve the precision of our ATE estimators.
We examine Figures 7, 8, and 9. The results indicate that Imaging tends to negatively impact students’ performance compared to Drawing in both proximal and distal periods. In contrast, Image Provided appears to enhance students’ performance in both periods. Additionally, while Drawing shows no significant difference compared to Image Provided in proximal performance, Image Provided demonstrates a positive effect on students’ distal performance compared to Drawing.
Finally, we examine the insights gained from Experiment 16, which investigated whether prompting students to reflect on their confusion before accessing a hint has impact on students’ performance. The findings are summarized in Table 5 and illustrated in Figure 10.
Table 5: Experiment 16: Reflect vs. Regular Hint
Performance Method Estimate Std. Error $\mathbf{Pr(>|t|)}$
Proximal $\hat{\tau}_{t-test}$ 0.0055 0.0248 0.8237
$\hat{\tau}_{OLS}$ 0.0072 0.0174 0.6776
$\hat{\tau}_{LOOP}$ 0.0101 0.0171 0.5543
Distal $\hat{\tau}_{t-test}$0.0060 0.0191 0.7555
$\hat{\tau}_{OLS}$0.0083 0.0109 0.4435
$\hat{\tau}_{LOOP}$0.0069 0.0113 0.5428
  • Note: The smallest standard error within each performance group is in bold.
Figure 10: Experiment 16: Confidence Intervals
Similar to the findings from the previous two experiments, $\hat{\tau}_{LOOP}$ is the recommended estimator due to its combination of a small standard error and unbiased property. Additionally, Figure 10 suggests that prompting students to reflect appears to enhance their performance during the proximal period but negatively impacts their performance in the distal term.
By reviewing these three experiments, we observe the clear advantages of using $\hat{\tau}_{LOOP}$ to estimate ATE. This estimator is both precise and easy to apply, making it a versatile choice for other experiments involving comparative tutoring designs. One notable observation is that the distal performance estimates have smaller standard errors and narrower confidence intervals. At first glance, this might seem counterintuitive, as distal performance represents a long-term imputation. However, this phenomenon makes sense when we consider that proximal performance was included as a super covariate in the estimation of distal performance. Since students’ current performance is highly predictive of their future performance, this inclusion improves the precision of the distal performance estimates.
Furthermore, to gain a deeper understanding of why these interventions enhance or impair students’ performance during the proximal or distal periods, additional analyses are needed. These could involve controlling for additional confounding factors or conducting psychological analyses to uncover the underlying mechanisms. Nonetheless, LOOP can be widely applied to various support comparison scenarios and provides valuable insights that can inform the design of future educational experiments.
6 Limitations, Future Plans, & Conclusions
6.1 Limitations
In this paper, we utilized data from E-Trials to evaluate whether different tutoring strategies have a significant positive or negative impact on students’ proximal and distal performance, where LOOP estimator are recommended to estimate ATE due to its reduced standard error. However, there are certain limitations associated with the methods employed.
  1. 1.
    The ATE estimation approach used in this paper was specifically tailored for support comparison experiments conducted in E-TRIALS. It won’t work for single support or problem varied studies, or studies from a different CBLP.
  2. 2.
    For the LOOP estimator to achieve a small standard error, it is crucial that $\hat{m}_{i}$ serves as a reliable estimator of $m_{i}$. Ensuring their closeness poses a significant challenge that must be addressed.
  3. 3.
    The LOOP method cannot handle categorical variables with more than $53$ levels, as observed when including $class\_xid$ as a covariate to adjust for students’ baseline performance. Further revisions to the LOOP package are necessary to overcome this limitation.
6.2 Future Plans
The method and code discussed in this paper are specifically designed for support comparison studies. One future direction is to develop a similar framework for single support studies. However, for problem-varied studies, creating a generalized method is less feasible due to the flexibility these studies provide researchers in designing and conducting experiments. This autonomy makes it challenging to standardize the process, as each study may require unique adaptations.
Another future plan is to build a Shiny application to fully automate the data cleaning and modeling process. This application would allow users to upload their CSV files and select, from a predefined list, their desired outcome variable. By automating this process, the Shiny application could reduce time and effort required to conduct experiments.
6.3 Conclusion
Despite the significant increase in the use of computer-based learning platforms, researchers still lack adequate tools to address the diverse challenges students face in these environments. These challenges range from a one-size-fits-all instructional approach to the delivery of culturally insensitive feedback messages. Such issues are a testimony to the urgent need for tools that can enable researchers to identify and implement effective solutions.
In this paper, we introduce E-TRIALS, a tool designed to help researchers make data-driven decisions to improve student performance in CBLPs. E-TRIALS supports multiple study types, each with its own advantages and disadvantages, enabling researchers to conduct experiments tailored to their specific needs. Using data from three support comparison studies, we developed a generalizable approach that can be applied to any support comparison experiment within E-TRIALS. Our work demonstrates the potential of E-TRIALS to advance research in the realm of learning sciences and contribute to the development of more effective, inclusive, and adaptive CBLPs.
7 Acknowledgments
We would like to thank NSF (e.g., 2118725, 2118904, 1950683, 1917808, 1931523, 1940236, 1917713, 1903304, 1822830, 1759229, 1724889, 1636782, & 1535428), IES (e.g., R305N210049, R305D210031, R305A170137, R305A170243, R305A180401, & R305A120125), GAANN (e.g., P200A180088 & P200A150306), EIR (U411B190024 & S411B210024), ONR (N00014-18-1-2768), NHI (R44GM146483), and Schmidt Futures. None of the opinions expressed here are that of the funders.